# MAPO: Advancing Multilingual Reasoning through Multilingual Alignment-as-Preference Optimization

### 12 Jan 2024 v1

@@ Abstract
Though reasoning abilities are considered language-agnostic, existing LLMs exhibit inconsistent reasoning abilities across different languages, e.g., reasoning in a pivot language is superior to other languages due to the imbalance of multilingual training data. To enhance reasoning abilities in non-pivot languages, we propose an alignment-as-preference optimization framework. Specifically, we adopt an open-source translation model to estimate the consistency between answers in non-pivot and pivot languages. We further adopt the answer consistency as the preference for DPO or PPO thus optimizing the lesser reasoning. Experiments show that our method significantly improves the model’s multilingual reasoning, with better reasoning consistency across languages. Our framework achieved a $13.7\%$ accuracy improvement on out-of-domain datasets MSVAMP while preserving the competitive performance on MGSM. Moreover, we find that iterative DPO is helpful for further alignment and improvement of the model’s multilingual mathematical reasoning ability, further pushing the improvement to $16.7\%$  [^1] The project will be available at https://github.com/NJUNLP/MAPO.
MAPO: Advancing Multilingual Reasoning through Multilingual Alignment-as-Preference Optimization
Shuaijie She , Shujian Huangthanks: Corresponding author , Wei Zou , Wenhao Zhu Xiang Liu, Xiang Geng, Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University shesj@smail.nju.edu.cn, huangsj@nju.edu.cn {zouw,zhuwh,liuxiang,gx}@smail.nju.edu.cn, chenjj@nju.edu.cn
@@ 1 Introduction
The reasoning ability of large-scale models has attracted much attention at present. Existing work improves the reasoning ability of models by constructing a large number of English fine-tuned data with corresponding inference processes Yu et al. (2023). Though, we consider reasoning to be language agnostic, existing studies Chen et al. (2023) show that due to the imbalance of pre-training and fine-tuning data across languages, the mathematical reasoning ability of existing large-scale models also varies across languages, e.g., English is far superior to that of the other languages.
To address the issue of insufficient reasoning ability in other languages,  Chen et al., 2023 translates the English questions with corresponding reasoning processes into other languages for supervised fine-tuning (Ouyang et al., 2022, SFT), thus improving the multilingual reasoning of the model. However, SFT is restricted by insufficient high-quality and diverse references with detailed reasoning in other languages, where the out-of-domain performance of the fine-tuned model is also hard to guarantee Ouyang et al. (2022). Furthermore, SFT also suffers from potential translation errors and catastrophic forgetting. Studies like Huang et al. (2023); Qin et al. (2023) circumvents LLM training by designing a two-stage prompting, which first requires the model to translate input problems from other languages into English for solutions. However, this strategy requires LLMs with proficiency in both translation and reasoning, which is considered even harder.
Intuitively, the reasoning process shall remain constant across languages, thus we can leverage the superior reasoning abilities in a pivot language, such as English, by aligning the reasoning process of non-pivot languages to the pivot. That is, we encourage problem-solving in non-pivot language to reason similar to that in the pivot language. Therefore, we propose a multilingual alignment-as-preference optimization (MAPO) framework that can improve the model’s multilingual reasoning ability with better generalization without annotating multilingual reasoning processes.
@@
Specifically, our framework consists of two parts: preference estimation via multilingual alignment, and preference optimization. During preference estimation, we collect the reasoning processes of different languages to the same question. Instead of utilizing direct translations, we adopt a well-trained translation model to yield the generation probability of the answers in pivot language given answers in non-pivot languages. Since the higher generation probability indicates a more consistent reasoning aligned with the pivot language, the corresponding answers in the non-pivot language are favored and shall be promoted.
Finally, we adopt state-of-the-art preference optimization methods, PPO and DPO, to optimize the alignment as preference. Experiments show that MAPO effectively improves reasoning consistency across languages, i.e., improves the lesser reasoning ability in non-pivot language, achieving a 13.7% improvement in accuracy on out-of-domain datasets MVSVAMP. MAPO also retains the original reasoning ability while generalizing to out-of-domain tests compared to the SFT baseline. Moreover, MAPO can further improve the multilingual reasoning of the model by iterative DPO, outperforming all baselines including ChatGPT OpenAI (2022) across all three benchmarks.
2 Related Work
@@ 2.1 Multilingual Reasoning
A straightforward measurement to evaluate the efficacy of large language models (LLMs) lies in their proficiency in tackling complex reasoning, e.g., their performance in solving mathematical reasoning problems. Recent work Wei et al. (2023); Wang et al. (2022) has verified a substantial improvement when LLMs are guided through a step-by-step reasoning process, instead of conducting a direct answer. Efforts to refine prompting strategies Wang et al. (2022); Yao et al. (2023); Lee and Kim (2023) or to fine-tune models using datasets annotated with explicit reasoning steps Yu et al. (2023); Yue et al. (2023) have yielded promising outcomes. Some research has introduced mathematical reasoning datasets in the form of application problems, such as GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022), and SVAMP Patel et al. (2021).
To evaluate the multilingual reasoning capabilities of LLMs, Shi et al., 2022 proposes MGSM by manually translating 250 samples of the GSM8K test set from English into nine other languages. Subsequent researches focus on enhancing the multilingual reasoning of LLMs. Chen et al., 2023 translates the GSM8K training data into the same nine languages for supervised fine-tuning (SFT), which improves the model’s multilingual reasoning capabilities. However, SFT suffers from data scarcity and catastrophic interference. Its out-of-domain generalization is also hard to guarantee. Multilingual reasoning via LLMs remains an open challenge.
@@ 2.2 Preference Optimization
Though pretrained language models on vast training data achieve outstanding performance amongst various tasks, they still suffer from potentially harmful behaviors such as fabricating facts and offensive replies. RLHF Zheng et al. (2023); Bai et al. (2022b) further rectify these model behaviors by reinforcement learning via the preference of human feedback. RLHF introduces a reward model $r_{\theta}(x,y)$ that captures the preference nuance from the human feedback. Then, $r_{\theta}(x,y)$ scores arbitrary LLM outputs $y$ given input $x$ for iterative policy rectifications during proximal policy optimization (Schulman et al., 2017, PPO). The tuning is guided by Eq 1 to maximize the expected rewards of the LLM policy $\pi_{\theta}$ with the minimum deviation from the SFT policy:
@@ 3 Method
Figure 1: Illustration of our alignment framework. For brevity, we only show three sampling results and simplify the optimization processes of DPO and PPO.
We propose a framework for multilingual alignment-as-preference optimization (MAPO). Specifically, our framework includes two parts: alignment as preference estimation and preference optimization.
3.1 Alignment as Preference Estimation
Intuitively, reasoning ability is language-agnostic. However, LLM reasoning varies across different languages, where we consider a dominant language provides the upper bound in reasoning for lesser languages. Therefore, a straightforward approach is to align reasoning in non-pivot languages to a dominant pivot language. That is, we can rectify LLM’s output reasoning in non-pivot languages by favoring those answers more aligned to the pivot language.
To circumvent potential errors introduced by direct translation evaluation, we simply adopt a well-trained multilingual translation model to calculate the ‘alignment’ scores between answers in non-pivot and pivot languages as a preference. Since the translation model is optimized on a large scale parallel data in source and target languages $(\mathbf{Y},\mathbf{\bar{Y}})$ by maximizing conditional generation probability in Eq 3:
@@ 3.1 Alignment as Preference Estimation
with more aligned target language $\mathbf{\bar{Y}}$ to $\mathbf{Y}$, the higher the conditional generation probability $P(\bar{Y}|Y)$. Then, we input the answers in non-pivot languages and force-decode a corresponding answer in pivot language for the ‘alignment’ score. Correspondingly, answers with higher ‘alignment’ scores in non-pivot languages are preferred during the following preference optimizations.
3.2 Alignment as Preference Optimization
With ‘alignment’ scores as a preference, we adopt two state-of-the-art preference optimizations: PPO and DPO, to optimize the alignment scores of reasoning in non-pivot languages given that in a dominant pivot language. To make life easier, since English is dominant in both reasoning and translation, we adopt English as the pivot language in the rest of the paper.
@@ 3.2.1 Optimization with PPO
$\mathcal{L}_{\text{ppo}}=\mathbb{E}_{i}[P(\bar{Y}|Y_{i},x)-\gamma\log\frac{\pi_{\theta}(Y_{i}|x)}{\pi_{\text{SFT}}(Y_{i}|x)}]$ (4)
@@ 3.2.3 Iterative DPO
Recent research such as LLaMA2 Touvron et al. (2023) and Anthropic Bai et al. (2022a) suggests that preference data updated through multiple iterative rounds can help synchronize the reward model and the policy model, which is beneficial for the performance of preference optimization.
@@ 4 Experiment
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 33.7 42.9 46.2 45.6 46.2 48.9 50.4 50.7 50.9 46.2 50.5
MathOctopusP 13B 35.2 46.8 42.8 43.2 48.8 47.6 44.4 48.0 48.4 45.0 53.2
base model: MetaMathOctopusP 7B
Base Model 29.9 40.8 44.3 44.0 47.9 48.4 51.2 52.4 50.1 45.4 53.8
DPO 42.5 52.1 53.2 59.1 54.2 55.7 56.5 58.8 58.5 54.5 61.2
base model: MathOctopusP 7B
Base Model 27.7 35.9 39.4 41.6 42.7 44.2 44.0 45.1 45.3 40.7 46.4
MultiLingual-RFT 37.9 46.4 46.4 49.6 50.8 50.4 50.7 51.6 53.4 48.6 49.4
PPO LoRA 38.9 47.5 47.1 51.0 51.7 51.1 50.3 51.6 51.4 49.0 52.7
DPO LoRA 39.4 44.7 48.9 50.8 53.7 53.2 53.3 54.2 55.7 50.4 56.5
DPO 42.0 53.2 52.7 54.7 56.4 56.9 55.7 58.5 59.3 54.4 59.6
DPO Iter1 45.3 53.9 53.8 56.8 58.1 56.6 58.7 59.1 58.5 55.6 60.0
DPO Iter2 48.8 55.2 56.0 60.3 58.8 58.3 58.1 59.7 60.8 57.3 58.4
Table 1: Model Performances on MSVAMP test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.
@@
Figure 3: ARC across nine languages on MSVAMP. Higher ARC indicate higher answer consistency.
4.1 Experiment Setup22footnotemark: 2
33footnotetext: More training details can refer to Appendix A.1
Base Models: Our experiments includes two base models: MathOctopusP-7B and MetaMathOctopusP-7B.  Chen et al. (2023) propose GSM8KInstruct by translating GSM8K Cobbe et al. (2021) to nine non-English languages. Based on the dataset, Chen et al. (2023) fine-tune the LLaMA2-7B Touvron et al. (2023), resulting in . To verify the effectiveness of our method in aligning models with stronger English reasoning abilities, we choose MetaMath-7B Yu et al. (2023) and continue to train it on GSM8KInstruct as another base model.
Alignment Datasets: NumGLUE Mishra et al. (2022) is a multi-task arithmetic reasoning benchmark. To obtain the mathematical world problems, we selected tasks 1, 4, and 8 from eight tasks as our dataset (NumGLUESub). Subsequently, we used ChatGPT to translate the questions into the same nine languages as GSM8KInstruct. NumGLUESub only has 1700 samples which is economical compared to GSM8KInstruct.
When constructing the preference dataset for DPO, we use MathOctopusP-7B as the sampling model and calculate the alignment feedback using NLLB-600M-distilled  [^4] https://huggingface.co/facebook/nllb-200-distilled-600M. We froze the parameters of both models and used them solely for inference. After that, the alignment dataset will be used for optimizing two base models respectively.
Evaluation Datasets:
  • MGSM (In domain) was gathered from a select portion of the GSM8K test set, after which native speaker rendered this subset into nine additional languages. Every linguistic variant is comprised of 250 individual test instances.
  • MNumGLUESub (In domain) We use ChatGPT to translate the test set of NumGLUESub to the same nine languages as MGSM which consists of 5530 samples in total.
  • MSVAMP (Out-of domain), obtained by translating SVAMP to the same nine language by Google Translate System and verified by three professional annotators. The dataset contains 1000 questions per language.
@@ 4.1 Experiment Setup22footnotemark: 2
  • PPL (PPL-based Alignment): We apply a translation model (NLLB-600M-distilled) to get the perplexity of the English answer given the non-English answer. Less perplexity indicates better alignment between the reasoning processes.
@@
Baselines: A strong baseline is the Base Model, which has acquired a degree of multilingual reasoning proficiency through fine-tuning on GSM8KInstruct. Moreover, motivated by rejection sampling techniques Yuan et al. (2023), we propose another strong baseline MultiLingual-RFT. We keep the reasoning solution which yield correct answers during sampling and then use these processes as labels to train the base model. To alleviate catastrophic forgetting, we fine-tune the model for only one epoch with lower learning rate (1e-5).
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 31.2 38.0 40.0 36.0 44.0 43.2 46.0 47.2 41.6 40.8 54.4
base model: MetaMathOctopusP
Base Model 28.8 34.0 36.4 34.0 42.8 42.0 42.0 44.8 41.2 38.4 52.4
DPO 28.8 34.8 37.6 41.2 47.2 44.0 46.8 43.6 38.0 40.2 53.6
base model: MathOctopusP
Base Model 29.2 33.6 36.4 35.2 39.2 38.8 44.8 42.4 43.2 38.1 52.0
MultiLingual-RFT 25.6 31.2 28.8 34.0 39.2 36.0 34.8 34.4 36.4 33.4 43.2
PPO LoRA 31.2 38.4 38.4 37.2 43.6 35.2 46.0 44.0 38.8 39.2 51.2
PPO LoRA 26.8 36.0 38.4 36.8 41.6 38.4 42.8 44.4 42.8 38.7 47.2
DPO 26.8 36.4 38.4 34.4 46.8 38.4 40.4 44.4 39.2 38.4 46.8
DPO Iter1 30.4 36.0 37.6 38.0 45.2 39.6 42.0 47.6 41.2 39.7 45.2
DPO Iter2 30.8 38.0 37.6 45.2 47.2 42.0 45.2 43.2 40.8 41.1 45.6
Table 2: Model Performances on MGSM test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 36.2 42.6 47.2 58.1 60.6 42.6 41.5 54.9 39.4 47.0 70.6
base model: MetaMathOctopusP
Base Model 24.5 35.2 36.2 44.1 45.8 39.2 32.6 45.2 36.7 37.7 52.5
DPO 37.3 42.0 41.6 51.6 54.8 44.4 48.2 52.4 45.2 46.4 64.0
base model: MathOctopusP
Base Model 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 36.0 46.9
MultiLingual-RFT 38.0 42.9 41.8 48.2 51.6 45.2 42.9 49.3 42.9 44.8 51.2
PPO LoRA 34.3 41.1 40.7 45.6 49.3 39.5 34.3 46.7 36.3 40.9 51.6
DPO LoRA 27.9 31.8 36.5 39.2 42.0 39.7 37.1 41.8 36.3 36.9 46.7
DPO 32.6 41.4 42.0 42.7 45.0 42.2 40.7 47.8 40.9 41.7 47.8
DPO Iter1 35.6 39.7 42.2 45.6 50.7 44.1 42.6 49.3 42.9 43.6 49.7
DPO Iter2 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 49.8 55.9
Table 3: Model Performances on MNumGLUESub test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.

### 22 Feb 2024 v2

# MAPO: Advancing Multilingual Reasoning through Multilingual-Alignment-as-Preference Optimization

@@ Abstract
Though reasoning abilities are considered language-agnostic, existing LLMs exhibit inconsistent reasoning abilities across different languages, e.g., reasoning in the dominant language like English is superior to other languages due to the imbalance of multilingual training data. To enhance reasoning abilities in non-dominant languages, we propose a Multilingual-Alignment-as-Preference Optimization framework (MAPO), aiming to align the reasoning processes in other languages with the dominant language. Specifically, we harness an off-the-shelf translation model for the consistency between answers in non-dominant and dominant languages, which we adopt as the preference for optimization, e.g., Direct Preference Optimization (DPO) or Proximal Policy Optimization (PPO). Experiments show that MAPO stably achieves significant improvements in the multilingual reasoning of various models on all three benchmarks (MSVAMP +16.2%, MGSM +6.1%, and MNumGLUESub +13.3%), with improved reasoning consistency across languages  [^1] The project will be available at https://github.com/NJUNLP/MAPO.
MAPO: Advancing Multilingual Reasoning through Multilingual-Alignment-as-Preference Optimization
Shuaijie She , Wei Zou , Shujian Huangthanks: Corresponding author , Wenhao Zhu Xiang Liu, Xiang Geng, Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University {shesj,zouw}@smail.nju.edu.cn, huangsj@nju.edu.cn {zhuwh,liuxiang,gx}@smail.nju.edu.cn, chenjj@nju.edu.cn
@@ 1 Introduction
The reasoning ability of large-scale models has attracted much attention Cobbe et al. (2021); Wang et al. (2022); Wei et al. (2023); Yu et al. (2023). Though we consider reasoning to be language-agnostic, existing studies Chen et al. (2023) show that due to the imbalance of pre-training and fine-tuning data across languages, the mathematical reasoning ability of existing large-scale models varies across languages, e.g., English, is far superior to that of the other languages.
To improve the reasoning ability in other languages,  Chen et al. (2023) translated the English reasoning processes into other languages for supervised fine-tuning (Ouyang et al., 2022, SFT). Although SFT brings in preliminary capability for multilingual reasoning, we argue that two problems hinder further improvement.
Firstly, the annotation for the reasoning process is expensive to obtain even for the dominant language, and the reasoning processes involve complex mathematical reasoning, which may result in translation errors Chen et al. (2023). As a result, the translated reasoning annotation for non-dominant languages is limited in both scale and quality. Without sufficient and diverse data, the results of supervised training are limited. It may also suffer from generalization issues in versatile task scenarios Zheng et al. (2023); Ouyang et al. (2022), indicating potential difficulty in adapting to the out-of-domain test set.
More importantly, although supervised training with translated reasoning processes improves the reasoning ability for almost all trained languages, this strategy only fills in the missing reasoning annotation for non-dominant languages that originate from the dominant language. Thus the inherent gap between dominant and non-dominant languages is hardly narrowed.
Different from existing attempts, we propose to use the reasoning ability of the dominant language as the director for improving non-dominant languages. As the reasoning process is critical for obtaining the correct result, the multilingual reasoning ability may be improved by encouraging the reasoning in non-dominant languages to be similar to that in the dominant language. Therefore, we propose a Multilingual-Alignment-as-Preference Optimization (MAPO) framework by aligning the reasoning process of non-dominant languages to the dominant. Notably, MAPO exploits the strong reasoning ability in the dominant language and requires no annotation for the reasoning process.
@@
More specifically, MAPO consists of two stages: preference estimation via multilingual alignment and preference optimization. During preference estimation, the reasoning processes to the same question are sampled from the LLM in both dominant and non-dominant languages. A well-trained, off-the-shelf translation model is employed to yield the translation probability between the reasoning in dominant and non-dominant languages. Since higher translation probability indicates a more consistent reasoning aligned with the dominant language, the corresponding reasoning in the non-dominant language is considered better and shall be promoted. During preference optimization, Proximal Policy Optimization (Schulman et al., 2017, PPO) and Direct Preference Optimization (Rafailov et al., 2023, DPO) are adopted to optimize the previously estimated preference. So the LLMs are trained to reason in the non-dominant languages as they do in the dominant language. We also conduct Iterative DPO for further preference optimization.
Experiments are conducted on three challenging multilingual reasoning test sets, namely MSVAMP Chen et al. (2023), MGSM Shi et al. (2022), and MNumGLUESub constructed from NumGLUE  Mishra et al. (2022), each covering 10 languages. MAPO achieves accuracy improvements of up to 16.2%, 6.1%, and 13.3% on the three benchmarks, respectively, reaching state-of-the-art performance on 7B models, thereby demonstrating the effectiveness of our method. Among them, the improvement on the out-of-domain dataset MSVAMP (+16.2%) shows that, by enhancing multilingual reasoning consistency, MAPO improves the multilingual reasoning capabilities of the model in a generalizable manner.
2 Preliminary
@@ 2.1 Multilingual Reasoning
A straightforward measurement to evaluate the efficacy of large language models (LLMs) lies in their proficiency in tackling complex reasoning, e.g., their performance in solving mathematical reasoning. Recent work Wei et al. (2023); Wang et al. (2022) has verified a substantial improvement when LLMs are guided through a step-by-step reasoning process, instead of conducting a direct answer.
Some research has introduced mathematical reasoning datasets in the form of application problems, such as GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022), and SVAMP Patel et al. (2021). To evaluate the multilingual reasoning capabilities of LLMs, Shi et al., 2022 propose MGSM by manually translating 250 samples of the GSM8K test set from English to other languages. Subsequent researches focus on enhancing the multilingual reasoning of LLMs. Chen et al., 2023 translate the GSM8K training data into other languages for supervised fine-tuning (SFT), which improves the model’s multilingual reasoning capabilities. However, SFT suffers from data scarcity and catastrophic forgetting. Its out-of-domain generalization is also hard to guarantee Zheng et al. (2023). Multilingual reasoning via LLMs remains an open challenge.
@@ 2.2 Preference Optimization
SFT maximizes the likelihood of annotated outputs and equips models with preliminary capabilities. However, models still exhibit various issues after SFT. Some researchers have further adjusted model behaviors and enhanced model capabilities through preference optimization.
RLHF Zheng et al. (2023); Bai et al. (2022b) further rectifies these model behaviors via reinforcement learning by preference. RLHF introduces a reward model $r_{\theta}(x,y)$ given input $x$ with its corresponding output $y$ that captures the preference nuance from the human feedback. Then, $r_{\theta}(x,y)$ scores arbitrary LLM outputs $y$ given input $x$ for iterative policy rectifications during proximal policy optimization (Schulman et al., 2017, PPO). The tuning is guided by Eq 1 to maximize the expected rewards of the LLM policy $\pi_{\theta}$ with the minimum deviation from the SFT policy:
@@
Figure 1: Illustration of our alignment framework. For brevity, we only show three sampling results and simplified optimization processes of DPO and PPO. The green and blue colors represent the same problem in Chinese and English, respectively. The white robots represent the original policy model and the colored robots represent the policy model with parameters updated through preference optimization.
@@ 3 Method
3.1 Preference Estimation
Intuitively, reasoning ability is language-agnostic. However, LLM reasoning varies across different languages, where we consider a dominant language provides the better reference in reasoning for lesser languages. Therefore, a straightforward approach is to align reasoning in non-dominant languages to a dominant language. That is, we can rectify LLM’s output reasoning in non-dominant languages by favoring those answers more aligned to the dominant language.
We simply adopt a well-trained multilingual translation model to calculate the ‘alignment’ scores between answers in non-dominant and dominant languages as a preference. Since the translation model is optimized on a large scale parallel data in source and target languages $(\mathbf{Y},\mathbf{\bar{Y}})$ by maximizing conditional generation probability in Eq 3:
@@ 3.1 Preference Estimation
with more aligned target language $\mathbf{\bar{Y}}$ to $\mathbf{Y}$, the higher the conditional generation probability $P(\mathbf{\bar{Y}}|Y)$. Then, we input the answers in non-dominant languages and force-decode a corresponding answer $\bar{Y}$ in dominant language for the ‘alignment’ score. Correspondingly, answers with higher ‘alignment’ scores in non-dominant languages are preferred during the following preference optimizations.
3.2 Preference Optimization
With ‘alignment’ scores as a preference, we adopt two state-of-the-art preference optimizations: PPO and DPO, to optimize the alignment scores of reasoning in non-dominant languages given that in a dominant language. To make life easier, since English is dominant in both reasoning and translation, we adopt English as the dominant language in the rest of the paper.
@@ 3.2.1 Optimization with PPO
$\mathcal{L}_{\text{PPO}}=\mathbb{E}_{i}[P(\bar{Y}|Y_{i},x)-\gamma\log\frac{\pi_{\theta}(Y_{i}|x)}{\pi_{\text{SFT}}(Y_{i}|x)}].$ (4)
@@ 3.2.2 Optimization with DPO
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y_{w}\mid x)}{\pi_{\text{ref}}(Y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(Y_{l}\mid x)}{\pi_{\text{ref}}(Y_{l}\mid x)}\right)\right].$ (5)
@@ 3.2.3 Iterative DPO
Recent research such as LLaMA2 Touvron et al. (2023) and Claude Bai et al. (2022a) suggests that preference data updated through multiple iterative rounds can is beneficial for the performance of preference optimization.
@@ 4 Experiment
4.1 Experiment Setup22footnotemark: 2
33footnotetext: More training details can refer to Appendix A
Base Models: Our experiments include two base models. MathOctopus-7B is fine-tuned from LLaMA2-7B with the multilingual reasoning processes MGSM8KInstruct [^4] GSM8K translated to nine non-English languages. Chen et al. (2023). MetaMathOctopus-7B is built from MetaMath-7B Yu et al. (2023), which has stronger English reasoning abilities, by fine-tuning with MGSM8KInstruct as well.
Preference Estimation: To obtain the mathematical problems for preference optimization, we selected tasks 1, 4, and 8 from the eight tasks in NumGLUE Mishra et al. (2022), which is a multi-task arithmetic reasoning benchmark, and translate the questions into the same nine languages as in MGSM8KInstruct, resulting the dataset MNumGLUESub. MNumGLUESub with only 1700 problems. When constructing the preference dataset for DPO, we use the corresponding base model for sampling and calculate the alignment feedback using NLLB-600M-distilled  [^5] https://huggingface.co/facebook/nllb-200-distilled-600M.
Preference Optimization : We experimented with both PPO and DPO. For simplicity, we report the results of the third iteration of DPO by default. For detailed results of PPO-LoRA and each round in Iterative DPO, please refer to Appendix B.
Dataset Type Number
MGSM In-Domain 2,500
MNumGLUESub In-Domain 5,530
MSVAMP Out-of-Domain 10,000
Table 1: Statistics of three benchmarks. “Type" indicates whether its corresponding training set was used during the training. “ Number” refers to the total number of samples in the test data.
Evaluation Datasets: We utilize three challenging benchmarks: MSVAMP, MGSM Shi et al. (2022) and MNumGLUESub (statistics are shown in Table 1). Among them, MSVAMP, serving as an Out-of-domain test set that does not participate in the training, is used to analyze the model’s robustness and generalization ability. MGSM is the testset corresponds to MGSM8KInstruct, on which the base models are trained. MNumGLUESub is the testset corresponds to the data for preference estimation and optimization.
Model Bn Th Sw Ja Zh Ru De Es Fr En Avg
GPT-3.5-Turbo 33.7 42.9 46.2 45.6 46.2 48.9 50.4 50.7 50.9 50.5 46.6
MathOctopus 13B 35.2 46.8 42.8 43.2 48.8 47.6 44.4 48.0 48.4 53.2 45.8
MAmmoTH ${}^{\dagger}$ 4.3 6.3 4.2 26.7 26.8 33.7 39.6 42.9 39.9 45.1 26.3
WizardMath ${}^{\dagger}$ 16.1 17.0 10.3 37.9 36.3 37.4 39.2 44.8 37.7 48.5 32.5
MetaMath 7B 14.8 17.7 14.8 51.8 54.4 59.4 59.6 63.2 62.0 64.8 46.2
MathOctopus 7B 27.7 35.9 39.4 41.6 42.7 44.2 44.0 45.1 45.3 46.4 41.2
+ m-RFT 37.9 46.4 46.4 49.6 50.8 50.4 50.7 51.6 53.4 49.4 48.7
+ MAPO-DPO(ours) 48.8 55.2 56.0 60.3 58.8 58.3 58.1 59.7 60.8 58.4 57.4
MetaMathOctopus 7B 36.1 47.5 49.4 51.3 54.5 53.6 56.6 60.0 57.2 64.2 53.0
+ m-RFT 44.8 54.2 56.2 58.4 57.7 56.2 59.2 59.3 57.8 63.1 56.7
+ MAPO-DPO(ours) 50.1 61.6 61.7 65.9 65.7 64.8 68.4 68.5 68.6 71.6 64.7
Table 2: Model Performances on MSVAMP test set. “Avg” represents the average performance on ten languages and bold text denotes the best results within the same size. Results marked with ${}^{\dagger}$ come from Chen et al., 2023.
@@ 4.1 Experiment Setup22footnotemark: 2
Figure 3: ACR across nine languages on MSVAMP. Higher ACR indicate higher answer consistency.
@@
  • PPL (PPL-based Alignment Score): We input the target non-English answer and apply NLLB-600M-distilled to get the perplexity of the given English answer. Less perplexity indicates better alignment between the reasoning processes.
@@
Baselines: The selected base models are already strong baselines, which have been fine-tuned on MGSM8KInstruct. Moreover, motivated by Rejection sampling Fine-Tuning (RFT) Yuan et al. (2023), we employ another strong baseline m-RFT, where the solutions which yield correct answers during sampling are used to further fine-tune the base model. To alleviate catastrophic forgetting, we fine-tune the model for only one epoch with a minor learning rate (1e-5). For comparison, we also incorporated other recent LLaMA2-base models: MAmmoTH Yue et al. (2023) is developed by utilizing diverse datasets for math instruction, while WizardMath Luo et al. (2023) employs Reinforcement Learning from Evol-Instruct (RLEIF). MetaMath 7B Yu et al. (2023) is fine-tuned from the strongest English reasoning dataset, MetaMathQA.